The demand function for product X is: Qd = 600 – 20Px + 0.02Y – 5Pr
The supply function is: Qs = –300 + 10Px
Where:
Qd = the quantity of X demanded
Qs = the quantity of X supplied
Px = the price of product X
Y = the average consumer income
Pr = the price of the related product R

If Y = R35 000 and Pr = R20, draw the precise demand curve for
product X
1
Expert's answer
2019-04-08T09:56:36-0400

If Y = R35 000 and Pr = R20, then Qd = 600 - 20Px + 0.02×35000 - 5×20 = 1200 - 20Px.

Px = 60 - Q/20, this function can be drawn as a demand curve, which is a downward-sloping line.